1. The cost of an adult ticket for a Saturday matinee at Turtle Cinema is $7. The cost of a student ticket is $5. For a particular Saturday matinee, the cinema sold 45 tickets and collected $265. How many student tickets were sold for the Saturday matinee?
2.What is the area of the quadrilateral with vertices at (1, 1), (4, 1), (7, 5) and (4, 5)?

Answer by jorel1380(3719)

x+y=45
7x+5y=265
5x+5y=225
2x=40
x=20
y=25
20 adult tickets and 25 student tickets were sold..